Hello AlexBriner,
I understand that Disk Enclosures 1, 2 and 3 appear to have issues on Loop B.
Before proceeding, please ensure Loop A is up and functional.
You may reseat the I/O Module B on Disk Enclosures 1, 2 and 3 - one at a time
Also reseat the cables connected to these I/O Modules. Check the LED status
Check the status in Command View EVA
If known-good spare I/O Modules and loop cables are available, you may swap the affected components one at a time and check the status.
For further diagnosis, please capture the event logs and configuration dump from this EVA and raise a support case with HPE.